homecomingchicken:

Solve the system of equations and choose the correct answer from the list of options. (4 points) x + y = −3 y = 2x + 2

snowflake0531:

Do you know the substitution method, elimination method, or graphing method?

homecomingchicken:

no im in 8th grade im kinda dumb can you just give me the answer

snowflake0531:

I can't give you the answer but i can bring you to the answer xd

homecomingchicken:

okay can you explain

snowflake0531:

We have x + y = −3 y = 2x + 2 and using the substitution method, we can substitute the second equation into the first making it x+ (2x+2) = -3 follow so far

homecomingchicken:

ok

snowflake0531:

So then solve for x 3x+2=-3 3x = -5 which means that \[x= - \frac {5} {3}\]

snowflake0531:

x+y=-3 -5/3 + y = -3 y=1 1/3